# Multichannel Conflict-Avoiding Codes for Expanded Scenarios

### v1

@@ I Introduction
I-A Single-Channel Conflict-Avoiding Codes
Let $\mathbb{Z}_{L}\triangleq\{0,1,\ldots,L-1\}$ denote the ring of residue modulo $L$. For $S\subseteq\mathbb{Z}_{L}$, let
$d^{*}(S)\triangleq\{a-b\,(\text{mod }L):\,a,b\in S,a\neq b\}$
denote the set of (nonzero) differences of $S$.
@@ I-A Single-Channel Conflict-Avoiding Codes
Let $L$ and $w$ be two positive integers with $L\geq w$. A conflict-avoiding code (CAC) $\mathcal{C}$ of length $L$ with weight $w$ is a collection of $w$-subsets, called codewords, of $\mathbb{Z}_{L}$ such that
$d^{*}(S)\cap d^{*}(S^{\prime})=\emptyset\quad\forall S,S^{\prime}\in\mathcal{C},S\neq S^{\prime}.$ (1)
The equation (1) is called the disjoint-difference-set property.
Let $\mathrm{CAC}(L,w)$ denote the class of all CACs of length $L$ with weight $w$. The maximum size of a code in $\mathrm{CAC}(L,w)$ is denoted by $K(L,w)$, i.e.,
$K(L,w)\triangleq\max\{|\mathcal{C}|:\,\mathcal{C}\in\mathrm{CAC}(L,w)\}.$
A code $\mathcal{C}\in\mathrm{CAC}(L,w)$ is called optimal if its code size achieves $K(L,w)$.
A $w$-subset $S\subseteq\mathbb{Z}_{L}$ is said to be equi-difference with generator $g\in\mathbb{Z}_{L}^{*}$ if $S$ is of the form $\{0,g,2g,\ldots,(w-1)g\}$. A CAC is called equi-difference if it entirely consists of equi-difference codewords. Let $\mathrm{CAC^{e}}(L,w)\subset\mathrm{CAC}(L,w)$ denote the class of all equi-difference codes and $K^{e}(L,w)$ be the maximum size among $\mathrm{CAC^{e}}(L,w)$. Obviously, $K^{e}(L,w)\leq K(L,w)$.
I-B Multichannel Conflict-Avoiding Codes
In a multiple access network with $M$ channels, a transmission pattern is represented as a zero-one $M\times L$ array, where $M$ denotes the number of channels and $L$ refers to the codeword length. The rows and columns of the array are further indexed by $\mathcal{I}_{M}\triangleq\{1,2,...,M\}$ and $\mathbb{Z}_{L}$, respectively. In this setting, a codeword with weight $w$ can be realized as a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$, in which each element is an ordered pair.
For two indices $i,j\in\mathcal{I}_{M}$, define the $(i,j)$ set of differences by
Note that, in the identical-index case, $D_{S}(i,i)=d^{*}(S_{i})$. $D_{S}$ is called the array of differences of the transmission pattern $S$. Two arrays of differences $D_{S}$ and $D_{S^{\prime}}$ are said to be disjoint, denoted by $D_{S}\cap D_{S^{\prime}}=\emptyset,$ if they are entry-wise disjoint.
Definition 2.
Let $M$, $L$ and $w$ be positive integers. A multichannel conflict-avoiding code (MC-CAC) $\mathcal{C}$ of length $L$ with weight $w$ in $M$ channels is a collection of $w$-subsets, also called codewords, of $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ such that
@@ I-B Multichannel Conflict-Avoiding Codes
The equation (3) is called the disjoint-difference-array property.
Let $\text{MC-CAC}(M,L,w)$ denote the class of all MC-CACs of length $L$ with weight $w$ in $M$ channels. The maximum size of a code in $\text{MC-CAC}(M,L,w)$ is denoted by $K(M,L,w)$, i.e.,
@@
I-C Contribution
@@
    Define the so-called exceptional codewords in MC-CACs, and derive a series of related properties by means of Kneser’s theorem. See Lemma 2 and Lemma 3 in Section II.
@@
    Obtain two classes of optimal single-channel CACs of lengths $p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$, where $p_{i}$ are primes and $w$ refers to the Hamming weight. See Theorems 7 and 10 in Section IV. The constructions of these codes can be extended to mixed-weight CACs.
@@
    Obtain optimal MC-CACs of length $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with two channels. See Theorem 13 in Section V. The construction of these codes can be extended to mixed-weight MC-CACs, which are defined for the first time in this paper.
@@
    Propose a construction of MC-CACs of length $(2\frac{w}{M}-1)L^{\prime}$, where $w$ and $M$ refer to the Hamming weight and the number of channels. See Theorem 14 in Section VI.
  5. 5.
    Derive a general upper bound on the maximum size of an MC-CAC, where the code length is with some restriction. This can be seen as a generalization of some previously-known results. See Theorem 4 in Section III.
  6. 6.
    Some applications to MC-CACs with at most one-packet per time slot property. See Theorems 17, 18 and 19 in Section VII-A.
@@ II Exceptional Codewords
In a (single-channel) CAC in $\mathrm{CAC}(L,w)$, a codeword $S$ is called exceptional if $|d^{*}(S)|<2w-2$ [1, 2, 5]. In this section, we will extend the concept of exceptional codewords to MC-CACs, together with some characterizations of structural properties.
@@ II-A Exceptional Pairs and Revisit of the Kenser’s Theorem
The following lists some results about the number of stabilizers of the difference set of an exceptional subset, which can be seen in [5, Corollary 1 and Lemma 1].
@@ II-B Exceptional Codewords in MC-CACs
Consider a subset $S\in\mathcal{I}_{M}\times\mathbb{Z}_{L}$. For $i\in\mathcal{I}_{M}$, define
$S_{i}\triangleq\{t:\,(i,t)\in S\}.$
Obviously, $|S|=\sum_{i\in\mathcal{I}_{M}}|S_{i}|$. We further define
$e_{S}\triangleq|\{i\in\mathcal{I}_{M}:\,S_{i}\neq\emptyset\}|$
@@ III General Upper Bounds on the Maximum Code Size
The upper bounds in Theorem 3 can be improved in the case when $M<w$, which is shown in the following theorem.
@@
The upper bounds stated in Theorem 3 and Theorem 4 remain valid when $M=1$. So, both they can be viewed as generalizations of Theorem 2.
@@ IV Single-Channel CACs
We first introduce the $p$-ary representation of a positive integer. given a positive integer $n$, let
@@
The following is a useful property of the $p$-ary representation of $c\in\mathbb{Z}_{p^{r}}^{*}$, which can also be found in [5, Proposition 3].
@@ IV-A Codeword Length $L=(w-1)p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
This subsection is devoted to generalizing the following result, which is due to [5, Theorem 6].
@@
We first introduce the following construction.
@@
$\widehat{Q}_{i}\triangleq\{(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ $\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,$
$g\in\mathcal{S}_{r_{i}}(Q_{i}),x_{t}\in\mathbb{Z}_{p_{t}^{r_{t}}}\text{ for }i<t\leq n\}.$
@@
$S_{a}=\{j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=0,1,2\ldots,w-1\}$
@@
Case 2: $1\leq j\leq w-2$. From (31) we have $j(1,g)=\pm k(1,h)$ in $\mathbb{Z}_{w-1}\times\mathbb{Z}_{p_{i}^{r_{i}}}$, where $1\leq k\leq w-2$. If $j(1,g)=k(1,h)$, it follows that $g=h$, where a contradiction occurs. If j(1,g)=k(1,h)j(1,g)=-k(1,h), we have j+k=w1j+k=w-1, and then get jg=(jw+1)hjg=(j-w+1)h (mod $p_{i}^{r_{i}}$). By considering the $p_{i}$-ary representation, both $jg$ and (jw+1)h(j-w+1)h are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Since $j\leq w-2<p_{i}$, we have $j=j_{0}$ and $(j-w+1)=(j-w+1)_{0}$ (mod $p_{i}$), namely, both $j$ and jw+1j-w+1 are in $\mathsf{L}_{0}$. It follows from Proposition 3 that $g,h\in\mathsf{L}_{t}$. More precisely,
@@
Therefore, $g_{t},h_{t}\in Q_{i}$ by assumption. By Eq. (25) and Proposition 3,
@@
Case 2: $p_{1}\leq 2w-2$. Since $0\in d(S)$ for $S\in\mathcal{E}$, if follows from Proposition 1(ii) that $H_{S}\subseteq d(S)$. Then, $H_{S}\cap H_{S^{\prime}}=\{0\}$ for any two distinct $S,S^{\prime}\in\mathcal{E}$. Moreover, since $H_{S}$ is uniquely determined by its order due to Proposition 2, we have $\gcd(|H_{S}|,|H_{S^{\prime}}|)=1$. Now, assume $p_{t}<2w-1\leq p_{t+1}$ for some $1\leq t\leq n$. Since $|H_{S}|\leq 2w-2<p_{t+1}$, for $S\in\mathcal{E}$, the prime factors of $|H_{S}|$ are in $\{p_{1},\ldots,p_{t}\}$. This concludes that $|\mathcal{E}|\leq t$ and $|H_{S}|$, $S\in\mathcal{E}$, are pairwise coprime. Therefore,
@@
The primes that satisfy the two conditions (26) and (27) are characterized in [5] with the help of Gauss’s lemma and laws of quadratic reciprocity, which lead to a series of optimal CACs of length $(w-1)p^{r}$ in [5, Corollary 2]. As an application of Theorem 7, the following corollary generalizes these optimal CACs.
@@
    $K(3L^{\prime},4)=(L^{\prime}-1)/2$ if $p_{i}\equiv\pm 1\ (\bmod{8})$ for all $i$,

### v2

@@ I Introduction
I-A Background
A conflict-avoiding code (CAC) [11] is a deterministic multiple access scheme in a single-channel system that operates without grant-based scheduling, time synchronization, or channel feedback. Despite this simplicity, a CAC ensures that each active user has at least one conflict-free transmission within a specified time duration by leveraging its strong cross-correlation properties. So, compared with probabilistic schemes, CACs are better suited for low-cost distributed networks under sporadic traffic that pursue ultra-high deadline-constrained reliability [16, 10]. Recently, with the aim of further increasing the maximal number of potential users that can be supported, CACs are extended to multichannel CACs (MC-CACs) in [1, 19] by extending their cross-correlation properties in two-dimensional (2D) domains. Note that protocol sequences [15] and rendezvous sequences [2] can also be regarded as variants of CACs designed to satisfy different performance guarantees, which involve different cross-correlation properties.
Consider a $K$-user multiple access system with $M$ orthogonal channels indexed by $\mathcal{I}_{M}\triangleq\{1,2,...,M\}$ and at most $w$ active users, where the time axis is partitioned into equal-length time slots, whose duration corresponds to the transmission time for one packet. The relative time offset (in unit of a time slot) of user $k$, denoted by $\tau_{k}$, can be an arbitrary nonnegative integer. Following [22], we assumed that all the time offsets are unknown to all the users, and can never be learned as the users receive no feedback from the channel. Tailored for this system, the adopted access scheme is defined as an MC-CAC (or a CAC when $M=1$), which is a collection of $K$ codewords, each of which consists of $w$ distinct elements in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ such that any two distinct codewords have at most one overlapping element for any relative shift in $\mathbb{Z}_{L}$ between them. Here, $\mathbb{Z}_{L}\triangleq\{0,1,\ldots,L-1\}$ denotes the ring of residue modulo $L$. Each user $k$ is assigned a unique codeword from this collection, and sends out a packet over channel $m$ at time slot $t+\tau_{k}$ if and only if user $k$ is active and the assigned codeword contains an element $(m,t+\tau_{k}$ (mod $L$)). As a result, each active user has at least one conflict-free transmission in a period of $L$ slots regardless of the integral relative time offsets. Note that the design of MC-CACs can be easily extended to that for the fully asynchronous model (i.e., $\tau_{k}$ may be any positive real number for each user kk[22, 32].
I-B Preliminaries
Consider a subset $S\in\mathcal{I}_{M}\times\mathbb{Z}_{L}$. For $m\in\mathcal{I}_{M}$, define
$S_{m}\triangleq\{t:\,(m,t)\in S\}.$
By viewing $S$ as a transmission pattern described in the previous subsection, $S_{m}$ represents the set of time slots during which packets are transmitted on the $m$-th channel. Obviously, $|S|=\sum_{m\in\mathcal{I}_{M}}|S_{m}|$. For two indices $i,j\in\mathcal{I}_{M}$, define the $(i,j)$ set of differences by
$D_{S}$ is called the array of differences of $S$. Two arrays of differences $D_{S}$ and $D_{S^{\prime}}$ are said to be disjoint, denoted by $D_{S}\cap D_{S^{\prime}}=\emptyset,$ if they are entry-wise disjoint.
The formal definition of MC-CACs is given below.
@@ I-B Preliminaries
Let $M$, $L$ and $w$ be positive integers. An MC-CAC $\mathcal{C}$ of weight $w$ in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ is a collection of $w$-subsets, also called codewords, of $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ such that
@@
The equation (2) is called the disjoint-difference-array property. Let $\text{MC-CAC}(M,L,w)$ denote the class of all such MC-CACs.
The maximum size of a code in $\text{MC-CAC}(M,L,w)$ is denoted by $K(M,L,w)$, i.e.,
@@
When $M=1$, an MC-CAC reduces to a (single-channel) CAC, where each codeword is simply a subset of $\mathbb{Z}_{L}$. Accordingly, MC-CAC(1,L,w)(1,L,w) and $K(1,L,w)$ reduce to CAC(L,w)(L,w) and $K(L,w)$, respectively. In this case, (2) can be simplified as
$d^{*}(S)\cap d^{*}(S^{\prime})=\emptyset\quad\forall S,S^{\prime}\in\mathcal{C},S\neq S^{\prime},$ (3)
by defining $d^{*}(S)\triangleq\{a-b\,(\text{mod }L):\,a,b\in S,a\neq b\}$, which is called the set of (nonzero) differences of $S$. The equation (3) is called the disjoint-difference-set property. Note that, in the identical-index case of (1), one has $D_{S}(i,i)=d^{*}(S_{i})$. A $w$-subset $S\subseteq\mathbb{Z}_{L}$ is said to be equi-difference with generator $g\in\mathbb{Z}_{L}^{*}$ if $S$ is of the form $\{0,g,2g,\ldots,(w-1)g\}$. A CAC is called equi-difference if it entirely consists of equi-difference codewords. Let $\mathrm{CAC^{e}}(L,w)\subset\mathrm{CAC}(L,w)$ denote the class of all equi-difference codes and $K^{e}(L,w)$ be the maximum size among $\mathrm{CAC^{e}}(L,w)$. Obviously, $K^{e}(L,w)\leq K(L,w)$.
I-C Related Works
A common design goal of MC-CACs (or CACs) is to maximize the code size, i.e., to find $K(M,L,w)$. When $M=1$, it was shown in [27] that $K(L,w)$ increases approximately with slope $(2w-2)^{-1}$ as a function of length $L$, for any fixed $w$. Meanwhile, an asymptotically tight upper bound of $K(L,w)$ was given in [28]. Based on some techniques in finite fields and some properties of cyclotomic cosets, [25, 33] provided some direct and recursive constructions of optimal CACs for general weights. Recently, some of the above results were generalized in [20] by the help of some techniques in additive combinatorics. For small $w$, the exact value of $K(L,3)$ is completely determined by [11, 8, 23, 4] for even $L$. As for odd length, $K(L,3)$ is determined for $L$ being some particular prime [11] and some composite number with particular factors [12, 3, 21, 24]. If only equi-difference codewords are concerned, $K^{e}(L,w)$ is obtained for some particular $L$ with $w=3$ in [31, 14] and with weight $w=4$ in [17, 13]. For more constructions and extensions of optimal CACs, one can refer to [26, 5, 6].
When it comes to MC-CACs, upper bounds of $K(M,L,w)$ of weight $w=3,4$ and $M\geq w$ were derived in [19], in which some cases were proven to be optimal by direct constructions based on some specific combinatorial designs. The exact value of $K(M=2,L,w=3)$ was determined in [18] for some particular length $L$. Under the assumption that $M\geq w$ and the prime factors of $L$ are all larger than or equal to 2w12w-1, [30] derived a general upper bound of $K(M,L,w)$ and obtained some optimal constructions. Meanwhile, a special class of MC-CACs under the restriction that each user has at most one-packet per time slot (AM-OPPTS) was considered in [30] as well.
I-D Contribution
To date, all known optimal constructions of MC-CACs in the literature [19, 18, 30] rely on techniques from combinatorial design theory. On the other hand, a variety of algebraic methods have been developed for constructing CACs (e.g., [25, 28, 20]). Motivated by this, in this paper we initiate the use of algebraic tools for constructing MC-CACs. Meanwhile, most existing results on MC-CACs are derived under the assumption $M\geq w$, which may be undesirable in resource-limited systems. So this paper focuses on the case $M<w$.
The considered length of CACs and MC-CACs in this paper is of the form $L=ap_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$, where $\gcd(a,p_{i})=1$, $1\leq i\leq n$, for some $a$ and primes $p_{1},\ldots,p_{n}$. As $\gcd(a,p_{i})=1$ for all $i$, we have $\mathbb{Z}_{L}\cong\mathbb{Z}_{a}\times\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$. A natural bijection between $\mathbb{Z}_{L}$ and $\mathbb{Z}_{a}\times\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$ is via the Chinese Remainder Theorem (CRT) [7], i.e., $\theta:\mathbb{Z}_{L}\to\mathbb{Z}_{a}\times\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$ by
$\theta(x)=(x\,(\bmod\,a),x\,(\bmod\,p_{1}^{r_{1}}),\ldots,,x\,(\bmod\,p_{n}^{r_{n}})).$ (4)
Therefore, a $w$-subset in $\mathbb{Z}_{L}$ can be simply put as a $w$-subset in $\mathbb{Z}_{a}\times\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$. The remaining algebraic tools will be introduced in Sections IVVI as needed.
In the classical study of CACs, upper bounds on $K(L,w)$ are typically derived by characterizing the structure and number of exceptional codewords, leading to tighter bounds [27, 20]. Although this idea has also appeared in the context of MC-CACs [19, 18], a systematic treatment is still lacking. In this paper, we introduce a general definition of exceptional codewords for MC-CACs. By applying Kneser’s theorem, we establish several structural properties and derive upper bounds on $K(M,L,w)$. These results not only refine those in [30], which only considered the case $M\geq w$, but also extend the underlying proof techniques.
@@ I-D Contribution
    Define the so-called exceptional codewords for MC-CACs, and derive a series of structural properties by means of Kneser’s theorem. These properties can be seen as generalizations of those of exceptional codewords for traditional CACs. See Lemmas 2 and 3.
@@
    Derive a general upper bound on the maximum code size of an MC-CAC, $K(M,L,w)$, for $M<w$, where the code length $L$ is under the assumption that the prime factors of $L$ are all larger than or equal to 2w12w-1. This not only refines the previously-known results in the literature, but also extend the underlying proof techniques.
@@
    Obtain two classes of optimal CACs of lengths $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and weight $w$, where $p_{i}$ are primes. Based on these two results, we further obtain optimal MC-CACs of length $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and weight $w$ for $M=2$ channels. We also derive the lower and upper bounds of $K(M,L,w)$ when $M$ divides $w$, where the code length is $L=(2\frac{w}{M}-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$.
@@
    Extend our results to AM-OPPTS MC-CACs. The construction of these codes are extended to mixed-weight MC-CACs, which are defined for the first time in this paper.
The rest of this paper is organized as follows. We define exceptional codewords for MC-CACs and provide some structural properties about them in Section II. A general upper bound of $K(M,L,w)$ is derived in Section III. Based on two new class of optimal CACs given in Section IV, we obtain a class of MC-CACs for two channels in Section V. Section VI investigates MC-CACs for $M$ channels when $M$ divides $w$. AM-OPPTS MC-CACs and mixed-weight MC-CACs are discussed in Section VII. Some concluding remarks including the growth rate of $K(M,L,w)$ are given in Section VIII.
@@ II Exceptional Codewords
In a CAC of weight $w$, a codeword $S$ is called exceptional if $|d^{*}(S)|<2w-2$ [28, 27, 20]. In this section, we will extend the concept of exceptional codewords to MC-CACs, together with some characterizations of structural properties.
@@ II-A Exceptional Pairs and Revisit of the Kenser’s Theorem
The following lists some results about the number of stabilizers of the difference set of an exceptional subset, which can be seen in [20, Corollary 1 and Lemma 1].
@@ II-B Exceptional Codewords in MC-CACs
For a subset $S\in\mathcal{I}_{M}\times\mathbb{Z}_{L}$, recall that $S_{m}$ collects time slots during which packets are transmitted on the $m$-th channel. We further define
$e_{S}\triangleq|\{m\in\mathcal{I}_{M}:\,S_{m}\neq\emptyset\}|$
@@ III General Upper Bounds on the Maximum Code Size
The upper bounds in Theorem 3 can be improved in the case when $M<w$, which is shown in the following theorem. For convenience, let $\mathbb{Z}_{L}^{*}=\mathbb{Z}_{L}\setminus\{0\}$ be the set of nonzeros in $\mathbb{Z}_{L}$.
@@
The upper bound stated in Theorem 3 and Theorem 4 remain valid when $M=1$. So, both they can be viewed as generalizations of Theorem 2.
@@ IV Single-Channel CACs
Firstly, we introduce the $p$-ary representation of a positive integer. Given a positive integer $n$, let
@@
The following is a useful property of the $p$-ary representation of $c\in\mathbb{Z}_{p^{r}}^{*}$, which can also be found in [20, Proposition 3].
@@ IV-A Codeword Length $L=(w-1)p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
This subsection is devoted to generalizing the following result, which is due to [20, Theorem 6].
@@
We first introduce the following direct construction.
@@
$\begin{split}\widehat{Q}_{i}\triangleq\{(0,\ldots,0,g,x_{i+1},\ldots,x_{n})&\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,\\ &g\in\mathcal{S}_{r_{i}}(Q_{i}),x_{t}\in\mathbb{Z}_{p_{t}^{r_{t}}}\text{ for }i<t\leq n\}.\end{split}$ (29)
@@
$S_{a}=\{j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=0,1,2\ldots,w-1\}$ (31)
@@
Case 2: $1\leq j\leq w-2$. From (34) we have $j(1,g)=\pm k(1,h)$ in $\mathbb{Z}_{w-1}\times\mathbb{Z}_{p_{i}^{r_{i}}}$, where $1\leq k\leq w-2$. If $j(1,g)=k(1,h)$, it follows that $j=k$ and hence $g=h$, where a contradiction occurs. If j(1,g)=k(1,h)j(1,g)=-k(1,h), we have j+k=w1j+k=w-1, and then get jg=(jw+1)hjg=(j-w+1)h (mod $p_{i}^{r_{i}}$). By considering the $p_{i}$-ary representation, both $jg$ and (jw+1)h(j-w+1)h are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Since $j\leq w-2<p_{i}$, we have $j=j_{0}$ and $(j-w+1)=(j-w+1)_{0}$ (mod $p_{i}$), namely, both $j$ and jw+1j-w+1 are in $\mathsf{L}_{0}$. It follows from Proposition 3 that $g,h\in\mathsf{L}_{t}$. More precisely,
@@
Therefore, $g_{t},h_{t}\in Q_{i}$ by assumption. By (26) and Proposition 3,
@@
Example 1.
Let $w=4,n=2,p_{1}=7,$ and $p_{2}=23$. The sets of quadratic residues modulo $7$ and $23$ are $Q_{1}=\{1,2,4\}$ and $Q_{2}=\{1,2,3,4,6,8,9,12,13,16,18\}$, respectively. Obviously, $\left(\frac{-1}{7}\right)=\left(\frac{1}{7}\right)\left(\frac{-2}{7}\right)=-1$ and $\left(\frac{-1}{23}\right)=\left(\frac{1}{23}\right)\left(\frac{-2}{23}\right)=-1$. By Theorem 6, we have an equi-difference CAC of length $L=3\cdot 7^{r}\cdot 23^{s}$ and weight $4$ with $(7^{r}\cdot 23^{s}-1)/2$ codewords, for any positive integers $r$ and $s$. When $r=s=1$, we have $\widehat{Q}_{1}=\{(g,x)\in\mathbb{Z}_{7}\times\mathbb{Z}_{23}:\,g\in Q_{1},x\in\mathbb{Z}_{23}\}$ and $\widehat{Q}_{2}=\{(0,g)\in\mathbb{Z}_{7}\times\mathbb{Z}_{23}:\,g\in Q_{2}\}$. The obtained code in $\mathrm{CAC^{e}}(483,4)$ has generators in $\{\theta^{-1}(1,a):\,a\in\widehat{Q}_{1}\cup\widehat{Q}_{2}\}$, where elements $(1,a)$ are considered in $\mathbb{Z}_{3}\times\mathbb{Z}_{7}\times\mathbb{Z}_{23}$ and the bijection $\theta:\mathbb{Z}_{483}\to\mathbb{Z}_{3}\times\mathbb{Z}_{7}\times\mathbb{Z}_{23}$ is given in (4). The corresponding generators are listed as follows.
$\{\theta^{-1}(1,1,x):\,x\in\mathbb{Z}_{23}\}$ $=\{1,22,43,64,85,106,127,148,169,190,211,232,$
$\qquad\!253,274,295,316,337,379,385,400,421,442,463\}$
$\{\theta^{-1}(1,2,x):\,x\in\mathbb{Z}_{23}\}$ $=\{16,37,58,79,100,121,142,163,184,205,226,247,$
$\qquad\!268,289,310,331,352,373,394,415,436,457,478\}$
$\{\theta^{-1}(1,4,x):\,x\in\mathbb{Z}_{23}\}$ $=\{4,25,46,67,88,109,130,151,172,193,214,235,$
$\qquad\!256,277,298,319,340,361,382,403,424,445,466\}$
$\{\theta^{-1}(1,0,g):\,g\in Q_{2}\}$ $=\{49,133,154,196,232,238,259,280,301,427,469\}.$
We now show that the construction described in Theorem 6 is optimal.
@@
Case 2: $p_{1}\leq 2w-2$. Since $0\in d(S)$ for $S\in\mathcal{E}$, if follows from Proposition 1(ii) that $H_{S}\subseteq d(S)$. Then, $H_{S}\cap H_{S^{\prime}}=\{0\}$ for any two distinct $S,S^{\prime}\in\mathcal{E}$. Moreover, $\gcd(|H_{S}|,|H_{S^{\prime}}|)=1$. Assume on the contrary that $\gcd(|H_{S}|,|H_{S^{\prime}}|)=d>1$. Since $L/d$ is a multiple of $L/|H_{S}|$ (also a multiple of $L/|H_{S^{\prime}}|$), it follows from Proposition 2 that $L/d\in H_{S}\cap H_{S^{\prime}}$, a contradiction. Now, assume $p_{t}<2w-1\leq p_{t+1}$ for some $1\leq t\leq n$. Since $|H_{S}|\leq 2w-2<p_{t+1}$, for $S\in\mathcal{E}$, the prime factors of $|H_{S}|$ are in $\{p_{1},\ldots,p_{t}\}$. This concludes that $|\mathcal{E}|\leq t$ and $|H_{S}|$, $S\in\mathcal{E}$, are pairwise coprime. Therefore,
@@
The primes that satisfy the two conditions (27) and (28) are characterized in [20] with the help of Gauss’s lemma and laws of quadratic reciprocity, which lead to a series of optimal CACs of length $(w-1)p^{r}$ in [20, Corollary 2]. As an application of Theorem 7, the following corollary generalizes these optimal CACs.
@@
    $K(3L^{\prime},4)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1\ (\bmod\ 8)$ for all $i$,